# disposition

### v2

@@ Abstract
Abstract
Measuring dependence between two events, or equivalently between two binary random variables, amounts to expressing the dependence structure inherent in a $2\times 2$ contingency table in a real number between $-1$ and 1. Countless such dependence measures exist, but there is little theoretical guidance on how they compare and on their advantages and shortcomings. Thus, practitioners might be overwhelmed by the problem of choosing a suitable measure. We provide a set of natural desirable properties that a proper dependence measure should fulfill. We show that Yule’s ${Q}$ and the little-known Cole coefficient are proper, while the most widely-used measures, the phi coefficient and all contingency coefficients, are improper. They have a severe attainability problem, that is, even under perfect dependence they can be very far away from $-1$ and $1$, and often differ substantially from the proper measures in that they understate strength of dependence. The structural reason is that these are measures for equality of events rather than of dependence. We derive the (in some instances non-standard) limiting distributions of the measures and illustrate how asymptotically valid confidence intervals can be constructed. In a case study on drug consumption we demonstrate how misleading conclusions may arise from the use of improper dependence measures.
@@ 1 Introduction
Analyzing dependence between two random variables is a fundamental task in statistics. The arguably simplest setting is the one of two binary random variables or, equivalently, two events, which is fully characterized by two marginal and a joint probability, and is conveniently represented in a $2\times 2$ contingency table. It is often useful to quantify the dependence structure inherent in the contingency table with a dependence measure, that is, a number lying between $-1$ and 1 and indicating direction of dependence via its sign and strength of dependence via the closeness of its absolute value to 1. However, there exist many dependence measures for this task and hardly any theoretical guidance on the advantages and shortcomings of those measures and on their relationships. In this paper, we provide such a theoretical guidance based on a formal discussion of dependence concepts for events, an axiomatic approach to dependence measures and a theoretical analysis of existing measures and their relationships, leading to clear recommendations for statistical practice. We also develop statistical inference for the recommended measures.
@@ 1.1 Motivating Example
Estimated Dependence Measures measure value 90% CI Phi coefficient 0.23 $[0.16,0.30]$ Yule’s ${Q}$ 0.86 $[0.59,0.96]$ Cole’s coefficient ${C}$ 0.83 $[0.44,0.96]$
@@
While countless measures for this task have been proposed (see Warrens, (2019) for a comprehensive overview of such measures and their use in different disciplines), their theoretical properties and relationships have hardly been studied. Thus, it is difficult to pick a measure in practice and, if multiple measures are used, to understand why they often lead to different conclusions. Due to the lack of theoretical guidance, practitioners understandably are likely to choose the most popular measures. Table 2 contains the hits of a recent search for eight measures on Google and Google Scholar as a proxy for their popularity in practice and academia together with their estimated values on the smallpox data example of Table 1. The most popular measures can clearly be found on the left side, namely the phi coefficient (referred to as Matthews correlation coefficient in the machine learning literature (Matthews,, 1975)) and contingency coefficients, of which Cramér’s V is the most well-known representative. [^1] Note that contingency coefficients are applicable to general contingency tables. Thus, the hits may overstate the importance of Cramér’s V for the specific $2\times 2$ case. Those measures take the rather small value of 0.23, suggesting weak positive dependence between the two events. In contrast, the measures on the right side lie between 0.57 and 0.83, indicating a rather strong dependence and being more in line with the impression one gets when looking at the contingency table. We demonstrate in this article that there is a reason for this impression and that indeed the measures on the right-hand side are theoretically superior to the ones on the left-hand side. The former, even though less widely-used, fulfill a set of crucial properties of dependence measures and thus give a sensible assessment of dependence.
Improper Dependence Measures measure value Google Scholar Cramér’s V $0.23$ $236{,}000$ $26{,}400$ Matthews cor. $0.23$ $202{,}000$ $34{,}400$ Phi coef. $0.23$ $94{,}200$ $24{,}500$ Pearson cont. coef. $0.23$ $4{,}650$ $880$
Proper Dependence Measures measure value Google Scholar Tetrach. cor. $0.61$ $34{,}400$ $9{,}490$ Yule’s ${Q}$ $0.86$ $18{,}200$ $2{,}120$ Yule’s ${Y}$ $0.57$ $4{,}530$ $429$ Cole’s coef. $0.83$ $460$ $131$
Table 2: Estimates of some improper (left) and proper (right) measures of dependence for the smallpox data example and Google and Google Scholar hits of those measures (30th October 2024).
@@ 1.2 Plan of the Paper
After a discussion of the covariance in the case of binary random variables, which serves as a building block for most of the other measures, Section 4 treats the popular, but improper, measures. We show that the phi coefficient, which is just a special case of the classical Pearson correlation coefficient, lacks the fundamental property of attainability. That is, it does in general not take the values 1 and $-1$ under perfect positive and negative dependence. We further demonstrate that it may take very small (absolute) values in those cases and usually strongly understates strength of dependence, where the severity of this problem depends on the marginal event probabilities. The phi coefficient consequently cannot provide a reliable assessment of strength of dependence and the usual interpretation of values close to 0 indicating weak and absolute values close to 1 indicating strong dependence does not apply here. While the phi coefficient can hence not be regarded as a suitable dependence measure, we show that it is instead useful for a different task, namely measuring closeness to equality of two events. This is for example required in the evaluation of binary classifiers, where equality of classification and outcome characterizes a perfect classifier. We also discuss several popular contingency coefficients and show that for $2\times 2$ contingency tables they are simple functions of the phi coefficient and thus inherit its shortcomings.
@@
Section 5 is dedicated to proper dependence measures. Cole’s coefficient is a very natural measure, just normalizing the covariance with its values under perfect positive and negative dependence. Yule’s ${Q}$ uses a different normalization, not requiring a case distinction between positive and negative dependence. We also touch upon a whole class of measures that generalize ${Q}$, including Yule’s ${Y}$, and on tetrachoric correlation. A measure that falls out of the classical framework of dependence measures lying between $-1$ and 1, but is extremely popular and nicely interpretable, is the odds ratio, which is closely connected to Yule’s ${Q}$ as well. We show that it is proper too.
@@
Our focus for the rest of the paper then lies on the phi coefficient, Yule’s ${Q}$ and Cole’s coefficient as basically all the other measures discussed are closely related to the first two and share their properties. We first compare their values under different marginal event probabilities, which demonstrates that the phi coefficient differs in general quite strongly from the proper measures and sheds light on their relation as well.
@@
In Section 6 we derive the asymptotic distributions of their empirical counterparts $\widehat{{C}}_{n}$, $\widehat{{Q}}_{n}$ and $\widehat{\phi}_{n}$, under assumptions covering independent and identically distributed observations, but also time series data. $\widehat{{Q}}_{n}$ and $\widehat{\phi}_{n}$ are asymptotically normal, allowing for tests and confidence intervals in the classical way. For $\widehat{{C}}_{n}$ the limiting distribution is highly nonstandard and involves case distinctions with respect to the value of the true ${C}$ due to the case distinction used in its normalization. While testing is straightforward, we construct possibly conservative confidence intervals by a Bonferroni-corrected combination of inverted tests that are built on the different asymptotic distributions. For all three measures, we also derive the limiting distribution of the Fisher transformation, which leads to a better approximation than the classical limit distribution when the measures lie close to $-1$ or 1 and employ them for testing and confidence intervals. E.g., the right side of Table 1 shows $90\%$-confidence intervals based on the Fisher transform for the data example.
@@ 1.3 Literature
The most popular dependence measures for binary random variables were already introduced around the year 1900. According to Ekström, (2011), the phi coefficient was independently proposed by Pearson, (1900), Boas, (1909) and Yule, (1912). Yule’s ${Q}$ was introduced by Yule, (1900) and tetrachoric correlation in Pearson, (1900). Yule, (1912) is a foundational paper too, amongst others introducing Yule’s ${Y}$ as a variant of Yule’s ${Q}$. In the subsequent decades, several contingency coefficients have been proposed such as Pearson’s contingency coefficient (Pearson,, 1904), Cramér’s V (Cramér,, 1945) and Tschuprow’s T (Tschuprow,, 1925). Cole’s coefficient was introduced by the ecologist Cole, (1949) and remained largely unnoticed by the statistical literature.
@@
Asymptotic sampling variances for estimators of $\phi$ and ${Q}$ are already discussed in the original papers cited above, but restricted to the case of independent and identically distributed data. Inference for Cole’s coefficient has not been developed yet.
@@ 2.1 Setup
Consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and two events $A,B\in\mathcal{F}$ with $10<\mathbb{P}(A)<$ 1 and $10<\mathbb{P}(B)<$ 1. We are interested in measuring direction and strength of dependence between those two events, or equivalently, between the two dichotomous or binary random variables $X:=\mathds{1}_{A}$ and $Y:=\mathds{1}_{B}$, which take the value 1 if $A$ or $B$, respectively, happen and 0 if not. Denote the complements of $A$ and $B$ by $\overline{A}$ and $\overline{B}$. The joint distribution is fully characterized by three of the four joint probabilities $\mathbb{P}(A\cap B)$, $\mathbb{P}(A\cap\overline{B})$, $\mathbb{P}(\overline{A}\cap B)$ and $\mathbb{P}(\overline{A}\cap\overline{B})$ or one joint and the two marginal probabilities, e.g., $\mathbb{P}(A\cap B)$, $\mathbb{P}(A)$ and $\mathbb{P}(B)$. It can conveniently be represented in a $2\times 2$ contingency table as in Table 3.
@@
$A$ $\mathbb{P}(A\cap B)$ $\mathbb{P}(A\cap\overline{B})$ $\mathbb{P}(A)$
@@ 2.3 Perfect Dependence
We now introduce concepts of perfect positive and negative dependence between events. We are not aware of a discussion of this key issue in the literature, even though a thorough treatment of dependence measures is not possible without such concepts. Fortunately, they are very natural as well. From Definition 2.2 we know that for fixed event probabilities the dependence between $A$ and $B$ gets stronger in the positive direction if the intersection $A\cap B$ grows. The dependence gets maximal if $A$ and $B$ occur together as often as possible. This is the case if the intersection reaches its maximal size, that is, if $\mathbb{P}(A\setminus B)=\mathbb{P}(A\cap\overline{B})=0$ or $\mathbb{P}(B\setminus A)=0$. Essentially, this means that the smaller event is a subset of the larger and thus the intersection is equal to the smaller event. Conversely, the dependence gets strongest in the negative sense if $A$ and $B$ occur together as little as possible, that is, if their intersection reaches the minimal size, i.e., if the intersection of one set and the complement of the other reaches the maximal size.
@@
$A$ and $B$ are perfectly positively dependent if $\mathbb{P}(A\setminus B)=0$ or $\mathbb{P}(B\setminus A)=0$. They are perfectly negatively dependent if $\mathbb{P}(A\setminus\overline{B})=0$ or $\mathbb{P}(\overline{B}\setminus A)=0$. We call $A$ and $B$ perfectly dependent if they are perfectly positively or perfectly negatively dependent.
@@
    $A$ and $B$ are perfectly positively dependent if and only if $\mathbb{P}(A\cap B)=\min(\mathbb{P}(A),\mathbb{P}(B))$. They are perfectly negatively dependent if and only if $\mathbb{P}(A\cap B)=\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1)$. It further holds that $\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1)\leq\mathbb{P}(A\cap B)\leq\min(\mathbb{P}(A),\mathbb{P}(B))$.
@@
    $A$ and $B$ are perfectly positively dependent if and only if it holds that $P(A|B)=1$ or $P(B|A)=1$. They are perfectly negatively dependent if and only if it holds that $P(A|\overline{B})=1$ or $P(\overline{B}|A)=1$.
@@
    $A$ and $B$ are perfectly positively dependent if and only if it holds that $\mathbb{P}(A\cap\overline{B})=0$ or $\mathbb{P}(\overline{A}\cap B)=0$. They are perfectly negatively dependent if and only if it holds that $\mathbb{P}(A\cap B)=0$ or $\mathbb{P}(\overline{A}\cap\overline{B})=0$.
@@
    $A$ and $B$ are positively (negatively) dependent if and only if $A$ and $\overline{B}$ are negatively (positively) dependent.
@@
    $A$ and $B$ are stronger positively (negatively) dependent than $A^{*}$ and $B^{*}$ if and only if $A$ and $\overline{B}$ are stronger negatively (positively) dependent than $A^{*}$ and $\overline{B^{*}}$.
@@
    $A$ and $B$ are perfectly positively (negatively) dependent if and only if $A$ and $\overline{B}$ are perfectly negatively (positively) dependent.

### v3

@@ Abstract
Measuring dependence between two events, or equivalently between two binary random variables, amounts to expressing the dependence structure inherent in a $2\times 2$ contingency table in a real number between 1-1 and 1. Numerous such dependence measures exist, but there is little theoretical guidance on how they compare and on their advantages and shortcomings. Thus, practitioners might be overwhelmed by the problem of choosing a suitable measure. We provide a set of natural desirable properties that a proper dependence measure should fulfill. We show that Yule’s Q{Q} and the little-known Cole coefficient are proper, while the most widely-used measures, the phi coefficient and all contingency coefficients, are improper. They have a severe attainability problem, that is, even under perfect dependence they can be very far away from 1-1 and $1$, and often differ substantially from the proper measures in that they understate strength of dependence. The structural reason is that these are measures for equality of events rather than of dependence. We derive the (in some instances non-standard) limiting distributions of the measures and illustrate how asymptotically valid confidence intervals can be constructed. In a case study on drug consumption we demonstrate how misleading conclusions may arise from the use of improper dependence measures.
@@ 1 Introduction
Analyzing dependence between two random variables is a fundamental task in statistics. The arguably simplest setting is the one of two binary random variables or, equivalently, two events, which is fully characterized by two marginal and a joint probability, and is conveniently represented in a $2\times 2$ contingency table. It is often useful to quantify the dependence structure inherent in the contingency table with a dependence measure, that is, a number lying between 1-1 and 1 and indicating direction of dependence via its sign and strength of dependence via the closeness of its absolute value to 1. However, there exist many dependence measures for this task and hardly any theoretical guidance on the advantages and shortcomings of those measures and on their relationships. In this paper, we provide such a theoretical guidance based on a formal discussion of dependence concepts for events, an axiomatic approach to dependence measures and a theoretical analysis of existing measures and their relationships, leading to clear recommendations for statistical practice. We also develop statistical inference for the recommended measures.
@@ 1.1 Motivating Example
Estimated Dependence Measures measure value 90% CI Phi coefficient 0.23 $[0.16,0.30]$ Yule’s Q{Q} 0.86 $[0.59,0.96]$ Cole’s coefficient C{C} 0.83 $[0.44,0.96]$
@@
While numerous measures for this task have been proposed (see Warrens, (2019) for a comprehensive overview of such measures and their use in different disciplines), their theoretical properties and relationships have hardly been studied. Thus, it is difficult to pick a measure in practice and, if multiple measures are used, to understand why they often lead to different conclusions. Due to the lack of theoretical guidance, practitioners understandably are likely to choose the most popular measures. Table 2 contains the hits of a recent search for eight measures on Google and Google Scholar as a proxy for their popularity in practice and academia together with their estimated values on the smallpox data example of Table 1. The most popular measures can clearly be found on the left side, namely the phi coefficient (referred to as Matthews correlation coefficient in the machine learning literature (Matthews,, 1975)) and contingency coefficients, of which Cramér’s V is the most well-known representative. [^1] Note that contingency coefficients are applicable to general contingency tables. Thus, the hits may overstate the importance of Cramér’s V for the specific $2\times 2$ case. Those measures take the rather small value of 0.23, suggesting weak positive dependence between the two events. In contrast, the measures on the right side lie between 0.57 and 0.86, indicating a rather strong dependence and being more in line with the impression one gets when looking at the contingency table. We demonstrate in this article that there is a reason for this impression and that indeed the measures on the right-hand side are theoretically superior to the ones on the left-hand side. The former, even though less widely-used, fulfill a set of crucial properties of dependence measures and thus give a sensible assessment of dependence.
Improper Dependence Measures measure value Google Scholar Cramér’s V $0.23$ 312,000312{,}000 29,20029{,}200 Matthews cor. $0.23$ 333,000333{,}000 42,20042{,}200 Phi coef. $0.23$ 122,000122{,}000 26,50026{,}500 Pearson cont. coef. $0.23$ 3,5903{,}590 $515$
Proper Dependence Measures measure value Google Scholar Tetrach. cor. $0.61$ 53,90053{,}900 10,10010{,}100 Yule’s Q{Q} $0.86$ 14,50014{,}500 2,1802{,}180 Yule’s Y{Y} $0.57$ 4,1104{,}110 $439$ Cole’s coef. $0.83$ $368$ $136$
Table 2: Estimates of some improper (left) and proper (right) measures of dependence for the smallpox data example and Google and Google Scholar hits of those measures (7th November 2025).
@@ 1.2 Plan of the Paper
After a discussion of the covariance in the case of binary random variables, which serves as a building block for most of the other measures, Section 4 treats the popular, but improper, measures. We show that the phi coefficient, which is just a special case of the classical Pearson correlation coefficient, lacks the fundamental property of attainability. That is, it does in general not take the values 1 and 1-1 under perfect positive and negative dependence. We further demonstrate that it may take very small (absolute) values in those cases and usually strongly understates strength of dependence, where the severity of this problem depends on the marginal event probabilities. The phi coefficient consequently cannot provide a reliable assessment of strength of dependence and the usual interpretation of values close to 0 indicating weak and absolute values close to 1 indicating strong dependence does not apply here. While the phi coefficient can hence not be regarded as a proper dependence measure, we show that it is instead useful for a different task, namely measuring closeness to equality of two events. This is for example required in the evaluation of binary classifiers, where equality of classification and outcome characterizes a perfect classifier. We also discuss several popular contingency coefficients and show that for $2\times 2$ contingency tables they are simple functions of the phi coefficient and thus inherit its shortcomings.
@@
Section 5 is dedicated to proper dependence measures. Cole’s coefficient is a very natural measure, just normalizing the covariance with its values under perfect positive and negative dependence. Yule’s Q{Q} uses a different normalization, not requiring a case distinction between positive and negative dependence. We also touch upon a whole class of measures that generalize Q{Q}, including Yule’s Y{Y}, and on tetrachoric correlation. A measure that falls out of the classical framework of dependence measures lying between 1-1 and 1, but is extremely popular and nicely interpretable, is the odds ratio, which is closely connected to Yule’s Q{Q} as well. We show that it is proper too, albeit with respect to a modified set of axioms that suits its scale and multiplicative nature (see Proposition A.3).
@@
Our focus for the rest of the paper then lies on the phi coefficient, Yule’s Q{Q} and Cole’s coefficient as basically all the other measures discussed are closely related to the first two and share their properties. We first compare their values under different marginal event probabilities, which demonstrates that the phi coefficient differs in general quite strongly from the proper measures and sheds light on their relation as well.
@@
In Section 6 we derive the asymptotic distributions of their empirical counterparts $\widehat{{C}}_{n}$, $\widehat{{Q}}_{n}$ and $\widehat{\phi}_{n}$, under assumptions covering independent and identically distributed observations, but also time series data. $\widehat{{Q}}_{n}$ and $\widehat{\phi}_{n}$ are asymptotically normal, allowing for tests and confidence intervals in the classical way. For $\widehat{{C}}_{n}$ the limiting distribution is nonstandard and involves case distinctions with respect to the value of the true C{C} due to the case distinction used in its normalization. While testing is straightforward, we construct possibly conservative confidence intervals by a Bonferroni-corrected combination of inverted tests that are built on the different asymptotic distributions. For all three measures, we also derive the limiting distribution of the Fisher transformation, which leads to a better approximation than the classical limit distribution when the measures lie close to 1-1 or 1 and employ them for testing and confidence intervals. E.g., the right side of Table 1 shows $90\%$-confidence intervals based on the Fisher transformation for the data example.
@@ 1.3 Literature
The most popular dependence measures for binary random variables were already introduced around the year 1900. According to Ekström, (2011), the phi coefficient was independently proposed by Pearson, (1900), Boas, (1909) and Yule, (1912). Yule’s Q{Q} was introduced by Yule, (1900) and tetrachoric correlation in Pearson, (1900). Yule, (1912) is a foundational paper too, amongst others introducing Yule’s Y{Y} as a variant of Yule’s Q{Q}. In the subsequent decades, several contingency coefficients have been proposed such as Pearson’s contingency coefficient (Pearson,, 1904), Cramér’s V (Cramér,, 1945) and Tschuprow’s T (Tschuprow,, 1925). Cole’s coefficient was introduced by the ecologist Cole, (1949) and remained largely unnoticed by the statistical literature.
@@
Asymptotic sampling variances for estimators of $\phi$ and Q{Q} are already discussed in the original papers cited above, but restricted to the case of independent and identically distributed data. Inference for Cole’s coefficient is developed in this paper for the first time.
@@ 2.1 Setup
Consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and two events $A,B\in\mathcal{F}$ with $0<\mathbb{P}(A)<1$ and $0<\mathbb{P}(B)<1$. We are interested in measuring direction and strength of dependence between those two events, or equivalently, between the two dichotomous or binary random variables $X:=\mathds{1}_{A}$ and $Y:=\mathds{1}_{B}$, which take the value 1 if $A$ or $B$, respectively, happen and 0 if not. Denote the complements of $A$ and $B$ by $\overline{A}$ and $\overline{B}$. The joint distribution is fully characterized by three of the four joint probabilities $\mathbb{P}(A\cap B)$, $\mathbb{P}(A\cap\overline{B})$, $\mathbb{P}(\overline{A}\cap B)$ and $\mathbb{P}(\overline{A}\cap\overline{B})$ or one joint and the two marginal probabilities, e.g., $\mathbb{P}(A\cap B)$, $\mathbb{P}(A)$ and $\mathbb{P}(B)$. It can conveniently be represented in a $2\times 2$ contingency table as in Table 3.
@@
AA $\mathbb{P}(A\cap B)$ $\mathbb{P}(A\cap\overline{B})$ $\mathbb{P}(A)$
@@ 2.3 Perfect Dependence
We now introduce concepts of perfect positive and negative dependence between events. We are not aware of a discussion of this key issue in the literature, even though a thorough treatment of dependence measures is not possible without such concepts. Fortunately, they are very natural as well. From Definition 2.2 we know that for fixed event probabilities the dependence between $A$ and $B$ gets stronger in the positive direction if the intersection $A\cap B$ grows. The dependence gets maximal if $A$ and $B$ occur together as often as possible. This is the case if the intersection reaches its maximal size, that is, if $\mathbb{P}(A\setminus B)=\mathbb{P}(A\cap\overline{B})=0$ or $\mathbb{P}(B\setminus A)=0$. Essentially, this means that the smaller event is a subset of the larger and thus the intersection is equal to the smaller event. Conversely, the dependence gets strongest in the negative sense if $A$ and $B$ occur together as little as possible, that is, if their intersection reaches the minimal size. This is equivalent to the intersection of one set and the complement of the other reaching the maximal size.
@@
AA and $B$ are perfectly positively dependent if $\mathbb{P}(A\setminus B)=0$ or $\mathbb{P}(B\setminus A)=0$. They are perfectly negatively dependent if $\mathbb{P}(A\setminus\overline{B})=0$ or $\mathbb{P}(\overline{B}\setminus A)=0$. We call $A$ and $B$ perfectly dependent if they are perfectly positively or perfectly negatively dependent.
@@
    AA and $B$ are perfectly positively dependent if and only if $\mathbb{P}(A\cap B)=\min(\mathbb{P}(A),\mathbb{P}(B))$. They are perfectly negatively dependent if and only if $\mathbb{P}(A\cap B)=\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1)$. It further holds that $\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1)\leq\mathbb{P}(A\cap B)\leq\min(\mathbb{P}(A),\mathbb{P}(B))$.
@@
    AA and $B$ are perfectly positively dependent if and only if it holds that $P(A|B)=1$ or $P(B|A)=1$. They are perfectly negatively dependent if and only if it holds that $P(A|\overline{B})=1$ or $P(\overline{B}|A)=1$.
@@
    AA and $B$ are perfectly positively dependent if and only if it holds that $\mathbb{P}(A\cap\overline{B})=0$ or $\mathbb{P}(\overline{A}\cap B)=0$. They are perfectly negatively dependent if and only if it holds that $\mathbb{P}(A\cap B)=0$ or $\mathbb{P}(\overline{A}\cap\overline{B})=0$.
@@
    AA and $B$ are positively (negatively) dependent if and only if $A$ and $\overline{B}$ are negatively (positively) dependent.
@@
    AA and $B$ are stronger positively (negatively) dependent than $A^{*}$ and $B^{*}$ if and only if $A$ and $\overline{B}$ are stronger negatively (positively) dependent than $A^{*}$ and $\overline{B^{*}}$.
@@
    AA and $B$ are perfectly positively (negatively) dependent if and only if $A$ and $\overline{B}$ are perfectly negatively (positively) dependent.